OpenStudy (anonymous):

Ok who can help with this problem Expand the expression ( 3 + 4(a+b))^2 and combine like terms. Simplify your answer as much as possible.

OpenStudy (anonymous):

\[(3+4(a+b)){2}=3^{2}+2\times 3\times 4(a+b)+16(a+b)^{2}\]

OpenStudy (anonymous):

Can you do the rest now?

OpenStudy (anonymous):

yeah thanks for your help

OpenStudy (anonymous):

I am very happy to help you

OpenStudy (anonymous):

ok what would you get for an answer I just want to check to see if I did the problem correctly

OpenStudy (anonymous):

\[16a ^{2}+16b ^{2}+24a+24b+32ab+9\]

OpenStudy (anonymous):

ok thanks i did make a mistake, but i saw what i did wrong thanks for your help

OpenStudy (anonymous):

It's my pleasure